OpenStudy (suwhitney):

Solve for x: ln(2x + 2) = 5. Give answer rounded off to 3 decimal places.

OpenStudy (phi):

make each side a power to the base "e" then use the rule \[ e^{\ln(x)} = x \] to simplify the left side. what do you get ?

OpenStudy (anonymous):

\[\ln(2x+2)=5<=> \ln(2x+2)=5*\ln(e)<=>\ln(2x+2)=\ln(e^{5})=>2x+2=e^5=>\] \[x=\frac{ e^5-2 }{ 2 }\]

OpenStudy (anonymous):

do you got it @suwhitney

OpenStudy (suwhitney):

Yes, I think so - thank yall :)